OpenStudy (anonymous):

Can someone double check my work? Multiply and Simplify Expression: (3s - 3)(4s - 5)

OpenStudy (anonymous):

\[12s^2 - 15s - 12s + 15\] = \[12s^2 - 27s + 15\]

OpenStudy (anonymous):

\[\huge \checkmark\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

oh wait, sorry \[\huge \checkmark\checkmark\]

OpenStudy (anonymous):

^.^

OpenStudy (e.mccormick):

So you double checked it... hehe. But you do that if it is right or wrong! =P Though, she was right.

OpenStudy (anonymous):

Go figure :)

OpenStudy (e.mccormick):

That woul be the lemniscate? The go figure?

OpenStudy (anonymous):

Lemniscate? \[\huge \color{blue}\infty\]

OpenStudy (anonymous):

For this next one it says find the product, but i did the same steps: (4a + 5)(4a - 5) \[16a^2 - 20a + 20a - 25\] \[16a^2 - 25\]

OpenStudy (anonymous):

Yup, @angelina22309 \[\huge \checkmark\] You'll find that \[\huge (a+b)(a-b)=a^2-b^2\]

OpenStudy (e.mccormick):

Yah, you go on the lemniscate and tell me when you find the end. Hehe. And Great angelina22309!

OpenStudy (anonymous):

oh the cleverness of me :)

OpenStudy (anonymous):

this one is a bit confusing, can you guys help? Multiply \[(x^4 + y^3)(x^4 - y^3) = x^8 -x^4y^3 + y^3x^4 - y^6\]

OpenStudy (e.mccormick):

Take a closer look at those middle two terms. Remember that xy=yx.

OpenStudy (e.mccormick):

Also, - means -1, so -xy=-1xy=-1yx

OpenStudy (anonymous):

so the middle terms \[-x^4x^3 +x^3y^4 = -x^7y^7 ?\] @e.mccormick

OpenStudy (e.mccormick):

Not quite. There is no multiplication there, so this is a negative added to a positive. Another term for tha is subtraction.

OpenStudy (anonymous):

Can you show me? I dont understand @e.mccormick

OpenStudy (e.mccormick):

Sure. You put this last:\[(x^4 + y^3)(x^4 - y^3) = x^8 -x^4y^3 + y^3x^4 - y^6\]I am saying change it to:\[(x^4 + y^3)(x^4 - y^3) = x^8 -y^3x^4 + y^3x^4 - y^6\]or\[(x^4 + y^3)(x^4 - y^3) = x^8 + y^3x^4-y^3x^4 - y^6\] Now, you have those center two terms: \(y^3x^4 -y^3x^4 \). What does that become?

OpenStudy (anonymous):

it would be: \[x^8 - y^6\]

OpenStudy (anonymous):

@e.mccormick

OpenStudy (e.mccormick):

You keep addressing them as if they are not a single multiplied term. Let me show you with what PeterPan pointed out: \[(a+b)(a-b)=a^2-b^2\] Well, if \(a=x^4\) and \(b=y^3\) then you get: \[(x^4+b^2)(x^4-b^2)=(x^4)^2-(y^3)^2=x^8-y^6\] What you have been missng is that those two center terms are the same, but their sign is different. Therefore they cancel.

OpenStudy (anonymous):

Once I wrote the answer I noticed that. Im not sure why I did not catch it when I started. @e.mccormick

OpenStudy (e.mccormick):

Yah, sometimes we just get stuck on an idea or forget some key fact and then we can't get it right. I worked for hours on smething trying to make it match the book answer because I thought mine was different. I never realized I had just brought things to the different side of the = sign, so my answer was both sides with opposite signs, which is the same value as what was in the book! And I could not see it!

OpenStudy (anonymous):

I'm going to post a fraction one (i always struggle with fractions) in another post. If you could help that would be great

OpenStudy (e.mccormick):

kk.